OpenStudy (anonymous):

HELP PLEASE!!!! Use a Sum or Difference Identity to evaluate sin 165°.

OpenStudy (anonymous):

well what two things when added would give you 165 degree's and ALSO that those two numbers would be degrees that you know on the unit circle?

OpenStudy (cwrw238):

sin 165 = sin (120 + 45) = sin120cos45 + cos120sin45 sin 120 = sqrt3/2, cos 120 = -1/2 sin 45 = cos45 = 1/sqrt2

OpenStudy (anonymous):

120 and 45

OpenStudy (cwrw238):

yea - thats what i used

OpenStudy (anonymous):

now do i just sub

OpenStudy (cwrw238):

yes sub those values in right hand side of the identity

OpenStudy (anonymous):

Yup, you can then just sub it in and carry out the calculations easy since they're on the unit circle. Here's a trig reference sheet i made too: http://learnix.net/wordpress/wp-content/uploads/Trig-Cheat-Sheet-1.4.pdf

OpenStudy (anonymous):

\[\sqrt{2}+\sqrt{6}/4\] \[\sqrt{6}-\sqrt{2}/4\] 1/2 -1

OpenStudy (anonymous):

which one is it??

OpenStudy (cwrw238):

i likt that reference sheet - really comprehensive

OpenStudy (anonymous):

@cwrw238 Thanks! :D

OpenStudy (anonymous):

@agentc0re or @cwrw238 which choice would u pick from what i posted??

OpenStudy (anonymous):

well i can't just give you the answer. :D What did you calculate?

OpenStudy (anonymous):

i got \[\sqrt{3}-1/2\sqrt{2}\] but thats nonw of my choices

OpenStudy (anonymous):

you're answer is correct. So they've done something, or did do something(simplify) to get the answers that they are giving you. That's a little annoying... Um, Multiply the answer by: \[\large \sqrt2 \over \sqrt2\]

OpenStudy (anonymous):

\[\sqrt{6}-\sqrt{2}/4\]

OpenStudy (anonymous):

now look at your answers. :D

OpenStudy (anonymous):

err answer choices. :D

OpenStudy (anonymous):

wait im wrong??

OpenStudy (anonymous):

No, you're right. it's also one of your answer choices. :D

OpenStudy (anonymous):

oh ok could u also help me with a few others??

OpenStudy (anonymous):

Did you download my reference sheet?

OpenStudy (anonymous):

yes y ?.

OpenStudy (anonymous):

@agentc0re r u there??

OpenStudy (anonymous):

i am.

OpenStudy (anonymous):

If sin =3/5 , find sin 2theta given that θ is in the 3rd quadrant.